# Fixed-Point RNNs: From Diagonal to Dense in a Few Iterations

### v1

@@ Abstract
Linear recurrent neural networks (RNNs) and state-space models (SSMs) such as Mamba have become promising alternatives to softmax-attention as sequence mixing layers in Transformer architectures. Current models, however, do not exhibit the full state-tracking expressivity of RNNs because they rely on channel-wise (i.e. diagonal) sequence mixing. In this paper, we propose to compute a dense linear RNN as the fixed-point of a parallelizable diagonal linear RNN in a single layer. We explore mechanisms to improve its memory and state-tracking abilities in practice, and achieve state-of-the-art results on the commonly used toy tasks $A_{5}$, $S_{5}$, copying, and modular arithmetics. We hope our results will open new avenues to more expressive and efficient sequence mixers.
Machine Learning, ICML
@@ 1 Introduction
State-space models (SSMs) and other new efficient recurrent token mixers are becoming a popular alternative to softmax attention in language modeling (Gu & Dao, 2023) as well as other applications such as vision (Liu et al., 2024), audio (Goel et al., 2022) and DNA processing (Nguyen et al., 2024). Inspired by linear input-controlled filtering, these models can be expressed as carefully parametrized linear recurrent neural networks (RNNs) with input-dependent, diagonal state transition: $h_{t}=\text{diag}(\mathbf{a}_{t})h_{t-1}+\mathbf{B}_{t}x_{t}$.
@@
Compared to classical RNNs such as LSTMs, the recurrence from the previous hidden state $h_{t-1}$ to the current $h_{t}$ is linear and its coefficient $\mathbf{a}_{t}$ does not depend on the hidden states. These factors allow SSMs such as Mamba to be computed through efficient parallel methods during training. At test time, they are faster than classical Transformers on long sequences due to their recurrent nature.
Figure 1: Sequence length generalization at training length 16 (pink) for state-tracking on $A_{5}$, with Transformer (brown) and LSTM (purple) as lower/upper bounds. Our Fixed-Point RNN (Bedouin) is trained at different maximum number of fixed-point iterations $\ell_{\text{max}}$: between $2$ (green) and $16$ (blue), or sampled from $\Gamma(4,1)$ with mean $4$ (gray). Increasing the number of fixed-point iterations turns the linear RNN from diagonal to dense in a few iterations.
@@
Though modern linear RNNs have shown promise in practice, recent theoretical studies suggest that using dense transition matrices (i.e. replacing $\text{diag}(\mathbf{a}_{t})$ with a dense $\mathbf{A}_{t}$) could present an opportunity to improve expressivity and unlock performance on challenging tasks. In particular, Cirone et al. (2024b) prove that dense linear SSMs/RNNs are endowed with the theoretical expressivity of classical non-linear RNNs such as LSTMs. As shown by Merrill et al. (2024), this is particularly useful in state-tracking applications where models are expected to maintain and extrapolate a complex state of the world. While state-tracking is naturally expressed in non-linear RNNs, it is known to be unavailable to channel-wise sequence mixers such as SSMs or Transformers (Merrill & Sabharwal, 2023). Finally, DeltaNet demonstrated that controlled, non-diagonal state updates significantly improve learning of compressed associative memories (Schlag et al., 2021a; Yang et al., 2024b).
Although linear recurrences are parallelizable across sequence length (Martin & Cundy, 2018), parallelizing dense RNNs efficiently is not trivial. Because channels and tokens need to be mixed jointly, structured transition matrices such as head-wise WY-representations (Yang et al., 2024b), are required for efficient implementations. But mixing only within heads reduces expressivity. In this paper, we present a general recipe to parametrize a large class of selective dense linear RNNs as fixed-points of diagonal linear RNNs.
Fixed-point RNNs.
In order to understand the relationship between dense linear RNNs and the fixed point of diagonal linear RNNs, consider the recurrent form of a dense RNN
$h^{*}_{t}=\mathbf{A}_{t}h^{*}_{t-1}+\mathbf{B}_{t}x_{t},$ (1)
where $\mathbf{A}_{t}\in\mathbb{R}^{d\times d}$ corresponds to the state transition matrix, $\mathbf{B}_{t}\in\mathbb{R}^{d\times d}$ corresponds to the input transition matrix, $h^{*}_{t}\in\mathbb{R}^{d}$ corresponds to the hidden state, and $x_{t}\in\mathbb{R}^{d}$ corresponds to the input for $t<T$ steps.
We parametrize the state transition matrix $\mathbf{A}_{t}$ as the product of a diagonal matrix $\Lambda_{t}\in\mathbb{R}^{d\times d}$, and a non-diagonal invertible mixing matrix $\mathbf{Q}_{t}\in\mathbb{R}^{d\times d}$
$h^{*}_{t}=\mathbf{Q}^{-1}_{t}\Lambda_{t}h^{*}_{t-1}+\mathbf{B}_{t}x_{t}.$ (2)
For the diagonal linear RNN $f_{\theta}(x,h)$, we use $\Lambda_{t}$ as the diagonal state transition matrix, and $\mathbf{Q}_{t}\mathbf{B}_{t}$ resp. $(\mathbf{I}-\mathbf{Q}_{t})$ as transition matrices for its two input sequences $x$ and $h$:
$f_{\theta}(x,h)_{t}=\Lambda_{t}f_{\theta}(x,h)_{t-1}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}.$ (3)
Assuming that the diagonal linear RNN has a fixed-point in depth $h^{*}=f_{\theta}(x,h^{*})\in\mathbb{R}^{T\times d}$, we can first reformulate
$h_{t}^{*}$ $=\Lambda_{t}h_{t-1}^{*}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{*}$
$\mathbf{Q}_{t}h_{t}^{*}$ $=\Lambda_{t}h_{t-1}^{*}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t},$
and than multiply both sides by $\mathbf{Q}_{t}^{-1}$ to obtain equation (2). This means that the states of the dense linear RNN are implicitly described by the fixed-point of $f_{\theta}(x,h)$ if it exists. Therefore, we can compute a dense linear RNN by finding the fixed-point of the corresponding diagonal linear RNN.
Motivated by this insight, we carefully parametrize the diagonal RNN and its channel mixer $\mathbf{Q}_{t}$ such that a fixed-point iteration converges towards the hidden states of the dense RNN. Intuitively, this introduces an iteration in depth of interleaved channel mixing with $\mathbf{Q}_{t}$ and sequence mixing with parallelizable linear recurrences (Martin & Cundy, 2018).
Contributions.
@@
    We develop a framework for Fixed-Point RNNs and establish its inherent compatibility with linear attention.
@@
    Fixed-Point RNNs trade parallelism for expressivity with the number of fixed point iterations (Figure 1).
@@
    Fixed-Point RNNs unite previously isolated capabilities of recurrent computation and memory (Figure 2).
Outline.
After discussing related work in Sec. 2, we introduce the fixed-point RNN framework in Sec. 3, and explore mechanisms for copying in Sec. 4. In Sec. 5, we apply the framework to Mamba and evaluate it in Sec. 6.
2 Related Work
Figure 2: (a) State-tracking on $A_{5}$ at sequence length $16$, and (b) character accuracy of copying at 2×2$\times$2 × sequence length generalization, trained on lengths $\in[5,50]$. Our single layer Fixed-Point RNN (Bedouin) with mixer reflections $r\in\{1,2,4\}$ is compared to baselines of increasing depth $\in\{1,2,4,6,8\}$. Bedouin is the only model capable of solving both the state-tracking and the copy task.
Beyond softmax attention.
The quadratic runtime complexity of Transformers motivated research on the linearization of its attention mechanism (Wang et al., 2020; Chen et al., 2021; Choromanski et al., 2020) – a technique that inevitably brings the sequence mixing mechanism closer to RNN-like processing (Katharopoulos et al., 2020; Schlag et al., 2021b). Recently, improvements on the long-range-arena benchmark (Tay et al., 2020) with state-space models (Gu et al., 2022; Smith et al., 2023) sparked a renewed interest in recurrent models (Gu & Dao, 2023; Sun et al., 2023; De et al., 2024; Qin et al., 2024; Peng et al., 2024; Yang et al., 2024a). New efficient token mixing strategies such as Mamba (Gu & Dao, 2023) showcase impressive results in language modeling (Waleffe et al., 2024) while offering linear runtime complexity. These models are fundamentally diagonal linear RNNs, which enables parallel algorithms such as parallel scans (Martin & Cundy, 2018) and fast linear attention based implementations (Yang et al., 2024b; Dao & Gu, 2024).
Expressivity of Diagonal vs. Dense RNNs.
Toy tasks.
Several works propose toy tasks to identify specific shortcomings of modern architectures. Specifically, Beck et al. (2024) use the Chomsky hierarchy to organize formal language tasks, of which a modular arithmetic task remains unsolved. With similar motivations, Merrill & Sabharwal (2023) introduce a set of word-problems for assessing state-tracking capabilities, among which the $A_{5}$ and $S_{5}$ tasks remain unsolved by Transformers and SSMs. Motivated by Transformers outperforming RNNs in memory capabilities, Jelassi et al. (2024) introduce a copying task as a fundamental benchmark for memory. We focus on these tasks to evaluate our Fixed-Point RNN framework.
Recurrence in Depth.
@@ 3 Fixed-Points as an RNN Layer
In this section, we first parametrize a diagonal linear RNN that is guaranteed to have an attracting fixed-point (Sec. 3.1). Then, we show that for the backward pass, it suffices to compute gradients only at the fixed-point $h^{*}$ itself (Sec. 3.2).
3.1 Computing the fixed point
Solving fixed-point equations such as $h=f_{\theta}(x,h)$, as needed from our discussion of (3), is perhaps one of the most well-studied problems in mathematics (Granas et al., 2003), often appearing in the fields of functional analysis and optimization. A common option in deep learning, used e.g. to find fixed points in deep equilibrium models (Bai et al., 2019), is Broyden’s method (Broyden, 1965). While much more efficient than standard optimization due to the lack of reliance on backward passes, Broyden’s method requires extra memory to keep track of the changes in the hidden state and is vulnerable to poorly conditioned curvatures (Bai et al., 2021; Martens, 2020).
However, focusing for the time being on the forward pass (see our discussion in Sec. 3.2 for the backward pass), a straightforward and yet effective method to compute the fixed point of an operator is simply to roll out the fixed point iteration. In the context of solving $h^{*}=f_{\theta}(x,h^{*})$, this corresponds to introducing an iteration in depth to (3). We introduce the superscript $\ell$ to the hidden state as $h_{t}^{\ell}$, where $t$ denotes the index of the current token (i.e., over the sequence dimension) and $\ell$ denotes the current iteration in depth (i.e., over the layer dimension) and compute $h_{t}^{\ell}$ from $h_{t}^{\ell-1}$ with $h_{t}^{0}=0$ as
$h_{t}^{\ell}=\Lambda_{t}h_{t-1}^{{\ell-1}}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{\ell-1}.$ (4)
The difficulty with such an iteration in time and depth is that, without any further constraint, it is prone to instabilities – due to the fact that fixed points might not exist or be unstable (Granas et al., 2003). While direct normalization techniques such as Layer Normalization might prevent from divergence (Beck et al., 2024), we here seek a more grounded alternative. Towards deriving a stable recurrence in depth, we make use of Banach’s fixed point theorem (Banach, 1922), that provides sufficient conditions for an operator to have an attracting fixed point:

### v2

# Fixed-Point RNNs: Interpolating from Diagonal to Dense

@@ Abstract
Linear recurrent neural networks (RNNs) and state-space models (SSMs) such as Mamba have become promising alternatives to softmax-attention as sequence mixing layers in Transformer architectures. Current models, however, do not exhibit the full state-tracking expressivity of RNNs because they rely on channel-wise (i.e. diagonal) sequence mixing. In this paper, we investigate parameterizations of a large class of dense linear RNNs as fixed-points of parallelizable diagonal linear RNNs. The resulting models can naturally trade expressivity for efficiency at a fixed number of parameters and achieve state-of-the-art results on the commonly used toy tasks $A_{5}$, $S_{5}$, copying, and modular arithmetics.
Figure 1: Sequence length generalization at training length 16 (pink) for state-tracking on $A_{5}$, with Transformer (brown) and LSTM (purple) as lower/upper bounds. Our Fixed-Point RNN (FP-Mamba) is trained at different maximum number of fixed-point iterations $\ell_{\text{max}}$: between $2$ (green) and $16$ (blue). Increasing the number of fixed-point iterations allows the linear RNN to interpolate from diagonal to dense in a few iterations.
@@ 1 Introduction
State-space models (SSMs) and other new efficient recurrent token mixers are becoming a popular alternative to softmax attention in language modeling (Gu & Dao, 2023) as well as in other applications such as vision (Liu et al., 2024) and DNA processing (Nguyen et al., 2024). Inspired by linear input-controlled filtering, these models can be expressed as carefully parametrized linear recurrent neural networks (RNNs) with input-dependent, diagonal state transition:
$h_{t}=\text{diag}(\mathbf{a}_{t})h_{t-1}+\mathbf{B}_{t}x_{t}$ (1)
@@
Compared to classical RNNs such as LSTMs (Hochreiter & Schmidhuber, 1997), in Eq. (1) the relation between the previous hidden state $h_{t-1}$ and the current $h_{t}$ is linear and its coefficient $\mathbf{a}_{t}$ does not depend on the hidden states. These choices allow SSMs such as Mamba (Gu & Dao, 2023) to be computed through efficient parallel methods during training. Furthermore, they are easier to optimize than classical RNNs, thanks to stable and efficient reparametrizations available for diagonal transitions (Orvieto et al., 2023; Zucchet & Orvieto, 2024) – techniques that are significantly more difficult to apply effectively in the classical setting (Arjovsky et al., 2016; Helfrich et al., 2018). At test time, they are faster than classical Transformers on long sequences due to their recurrent nature.
@@
Though modern linear RNNs have shown promise in practice, recent theoretical studies suggest that using dense, input-dependent transition matrices (i.e. replacing $\text{diag}(\mathbf{a}_{t})$ with a dense $\mathbf{A}_{t}$) could present an opportunity to improve expressivity and unlock performance on challenging tasks. In particular, Cirone et al. (2024b) prove that dense selective SSMs are endowed with the theoretical expressivity of classical non-linear RNNs such as LSTMs. As shown by Merrill et al. (2024) and Sarrof et al. (2024), such gained expressivity proves to be particularly useful in state-tracking applications where models are expected to maintain and extrapolate a complex state of the world. Since state-tracking is naturally expressed by non-linear RNNs but provably unavailable to channel-wise sequence mixers such as SSMs or Transformers, Merrill & Sabharwal (2023a) speculate on a fundamental tradeoff between parallelism and expressivity. This discussion sparked interest in non-diagonal recurrences and parallelizable architectures capable of state-tracking (Grazzi et al., 2024; Terzic et al., 2025; Schöne et al., 2025; Peng et al., 2025; Siems et al., 2025).
When designing new architectures involving dense selective yet linear state transitions of the form $h_{t}=\mathbf{A}_{t}h_{t-1}+\mathbf{B}_{t}x_{t}$, two fundamental concerns arise:
  1. 1.
    What should the parametric form for $\mathbf{A}_{t}$, as a function of the input be? How can we guarantee this parametrization induces a stable recurrence, like in standard [^1] Standard SSMs are diagonal and operate in polar coordinates, parametrizing directly the gap between eigenvalues and the stability threshold (Orvieto et al., 2023). This technique allows to increasing granularity near the identity, and to effectively normalize the forward pass (cf. $\sqrt{1-|\gamma|^{2}}$ term in Griffin (De et al., 2024)). SSMs?
  2. 2.
    How does a parametrization balance between expressivity and parallelism? Which assumptions on the structure of $\mathbf{A}_{t}$ enable efficient computation, and how do they interact with expressivity?
Perhaps the first approach tackling the above questions was DeltaNet (Schlag et al., 2021; Yang et al., 2024b) with a block-diagonal and orthogonal therefore, stable state transition structure, where each block is parametrized by a Householder matrix. The parallelizable algorithm, was then extended to include negative eigenvalues (Grazzi et al., 2024), gates (Yang et al., 2025), and most recently products of Householders (Siems et al., 2025). Such choices, leading to increased expressivity as exemplified by their state-tracking and length generalization capabilities, are motivated mainly by hardware considerations: Householder-based mixing can be implemented efficiently on GPUs as linear attention via WY-representations and the UT transform (Yang et al., 2024b).
Figure 2: (a) State-tracking on $A_{5}$ at sequence length $16$, and (b) character accuracy of copying at 2×2$\times$2 × sequence length generalization, trained on lengths $\in[5,50]$. Our single layer FP-Mamba with mixer reflections $r\in\{1,2,4\}$ is compared to baselines of increasing depth $\in\{1,2,4,6,8\}$. FP-Mamba is the only model capable of solving both the state-tracking and the copy task.
While the works above offer exciting practical strategies for boosting capabilities at a relatively low additional computational cost, they fall short in exploring the sea of intriguing options for dense transitions and hence, in thoroughly answering questions (1) and (2) above.
Unfortunately, this is not an easy task: although linear recurrences are theoretically parallelizable across sequence length (Martin & Cundy, 2018), parallelizing dense RNNs efficiently is not trivial due to increased memory I/O. These thoughts inspired us to change our viewpoint: instead of designing an algorithm which adds a fraction of non-diagonal processing to a model, here, we look for a strategy to navigate the parallelism tradeoff towards a truly dense object.
Motivated by the idea of designing a parallelizable general-purpose method to implement new dense RNN variations, in this paper we devise a new adaptive computation strategy which allows to interpolate between fast recurrent diagonal RNNs and dense recurrences with arbitrary preselected structure. Instead of parametrizing the dense RNN layer as an explicit function $h=F_{\theta}(x)$, we build on the literature of equilibrium/implicit models (Bai et al., 2019; Ghaoui et al., 2021) to parametrize it implicitly as a solution $h^{*}$ to a fixed-point equation $h=f_{\theta}(x,h)$ involving only a diagonal RNN. As described in Fig. 9(a), we solve for $h^{*}$ using a fixed-point iteration of diagonal RNN evaluations $f_{\theta}$.
A fundamental question some readers might rightfully ask, is the following: “what is the advantage of iterating a single layer in depth compared to depth-stacking multiple SSM, e.g. Mamba layers?” We claim one advantage comes from having access to the limiting dense object. As showcased by Fig. 2, this allows to adaptively provide the required expressivity for a fixed set of parameters without any a priori choice on the network size.
Summary.
@@
    We develop the framework of Fixed-Point RNNs to adaptively trade parallelism for expressivity using the number of fixed-point iterations (Fig. 1).
@@
    We achieve a stable parametrization of a dense RNN via a carefully designed diagonal RNN.
@@
    The framework allows for easy integration of both non-linear hidden state dependence and linear attention based matrix-valued formulations. This way, our FP-Mamba unites previously isolated capabilities of recurrent computation and memory (Fig. 2).
2 Background
Since their introduction (Rumelhart et al., 1986; Elman, 1990), RNNs have significantly contributed to the evolution of machine learning methods for sequential data (Hochreiter & Schmidhuber, 1997; Jaeger, 2001). But despite their theoretical promise of Turing-completeness (Siegelmann & Sontag, 1992), recurrent models fell out of fashion due to two significant challenges: they are inherently sequential, and notoriously difficult to train (Hochreiter et al., 2001; Pascanu et al., 2013). The recent advancements of linear RNNs (Gu & Dao, 2023) suggest a way forward to combine the scalability of Transformers (Vaswani et al., 2017) with the expressivity of classical RNNs (Cirone et al., 2024b). The key challenge here is the stable and efficient parametrization of a linear RNN layer with a time-varying recurrent transition matrix. In this paper we are exploring first steps towards this goal.
Dense Selective RNN.
Traditionally, RNNs are parametrized as either time-invariant, non-linear, or element-wise system. To the best of our knowledge, a time-variant, dense, and linear RNN parametrization has been of mild interest at best. To understand why, consider the general form
$F_{\theta}:x\mapsto h,$ $h_{t}=\mathbf{A}_{t}h_{t-1}+\mathbf{B}_{t}x_{t},$ (2)
where $\mathbf{A}_{t}\in\mathbb{R}^{d\times d}$ corresponds to the time-varying state transition matrix, $\mathbf{B}_{t}\in\mathbb{R}^{d\times d}$ is the input transformation matrix, $h_{t}\in\mathbb{R}^{d}$ denotes the hidden state, and $x_{t}\in\mathbb{R}^{d}$ is the input for $t<T$ steps. For a given sequence of $\mathbf{A}_{t}$, the complexity of a forward pass is $O(Td^{2})$ in memory and $O(T)$ sequential steps. Although such a linear RNN could also be computed in $O(\log T)$ sequential steps using a parallel scan algorithm (Martin & Cundy, 2018), this would require materializing matrix-matrix multiplications at cost $O(d^{3})$. An issue in both scenarios, however, is the parametrization of $\mathbf{A}_{t}$ as time-varying, i.e. input- or even hidden state-dependent matrices. In general, this requires a map $\mathcal{M}:d\mapsto d^{2}$ with potentially $d\times d^{2}$ parameters, and $O(Td^{3})$ time complexity. While structured dense matrix representations for $\mathbf{A}_{t}$ could potentially present a remedy, they come with additional challenges: (1) In order to guarantee expressivity, the $\mathbf{A}_{t}$ cannot be co-diagonalizable such as for example Toeplitz matrices (Cirone et al., 2024b). (2) In order to guarantee stability of the dynamical system, the spectral radius $\rho(\mathbf{A}_{t})$ needs to be less than, but still close to $1$ for long-range interactions (Orvieto et al., 2023). (3) The matrix structure needs to be closed under multiplications to enable parallel scans without having to materialize dense representations at $O(Td^{2})$ memory cost.
Related Works.
Improving the trainability of classical non-linear RNNs has a long history. For example, Arjovsky et al. (2016) and Helfrich et al. (2018) investigate parameterizations to stabilize their spectral radius with structured matrix representations, while Lim et al. (2024) and Gonzalez et al. (2024) propose iterative methods to parallelize their computation. In this work, however, we focus on stabilizing and parallelizing a time-variant, dense, linear RNN. Improving the limited expressivity of existing diagonal linear RNNs is the focus of a few recent works, e.g. by Grazzi et al. (2024) and Siems et al. (2025). In contrast, we investigate a wide class of structured parameterizations for dense RNNs where the additional cost is adaptively chosen depending on the task. In concurrent work, Schöne et al. (2025) propose an iterative method similar to ours, but as opposed to our carefully designed implicit dense RNN layer, they focus on scaling implicit causal models of existing multi-layer architectures on language. For a more extensive literature review, we refer the reader to App. A.
@@ 3 Fixed-Points as an RNN Layer
In this section, we introduce an implicit parameterization for a family of dense RNNs $F_{\theta}(x)$ which describes its output by a solution $h^{*}\in\mathbb{R}^{T\times d}$ to the fixed-point equation $h=f_{\theta}(x,h)$ (Sec. 3.1). Then, we discuss how to find the solution $h^{*}$ using fixed-point iterations (Sec. 3.2) and the algorithmic implications (Sec. 3.4) of the FP-RNNs framework in light of the challenges outlined in Sec. 2. Finally, we briefly touch on how to train an implicitly dense model $F_{\theta}(x)$ with gradient descent (Sec. 3.5).
3.1 From Explicit to Implicit Parameterization
We start by designing a diagonal RNN $f_{\theta}(x,h)$ such that the solution $h^{*}$ to its fixed-point equation $h=f_{\theta}(x,h)$ implicitly represents a dense RNN $h^{*}=F_{\theta}(x)$. Consider the factorized parametrization of $\mathbf{A}_{t}$ similar to the one introduced by Helfrich et al. (2018) for non-linear and time-invariant RNN:
$F_{\theta}:x\mapsto h^{*},$ $h^{*}_{t}=\mathbf{Q}^{-1}_{t}\mathbf{\Lambda}_{t}h^{*}_{t-1}+\mathbf{B}_{t}x_{t}.$ (3)
Separating $\mathbf{A}_{t}$ into a diagonal matrix $\mathbf{\Lambda}_{t}\in\mathbb{R}^{d\times d}$ and a non-diagonal invertible mixing matrix $\mathbf{Q}_{t}\in\mathbb{R}^{d\times d}$ allows to describe $h^{*}$ by only a diagonal transition $\mathbf{\Lambda}_{t}$ by reformulating Eq. 3 to
$h_{t}^{*}$ $=\mathbf{\Lambda}_{t}h_{t-1}^{*}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{*}.$ (4)
This means that the states $h^{*}=F_{\theta}(x)$ of the dense linear RNN can be implicitly described by the fixed-point $h^{*}=f_{\theta}(x,h^{*})$ of a corresponding diagonal linear RNN of the following form:
$f_{\theta}:(x,h)\mapsto h^{\prime},$ $h^{\prime}_{t}=\mathbf{\Lambda}_{t}h^{\prime}_{t-1}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}.$ (5)
In other words, if we could find the fixed-point $h^{*}=f_{\theta}(x,h^{*})\in\mathbb{R}^{T\times d}$ for the diagonal RNN defined in Eq. 5, then $h^{*}$ would describe the states of a corresponding dense RNN $h^{*}=F_{\theta}(x)$. Motivated by this insight, in Sec. 3.2 we carefully parametrize the diagonal RNN $f_{\theta}(x,h)$ and its channel mixer $\mathbf{Q}_{t}$ such that a computable fixed-point exists.
((a)) Framework
((b)) State Tracking on $A_{5}$
((c)) Number of Iterations $\ell^{*}$
Figure 3: (a) An overview of the proposed Fixed-Point RNN framework in Sec. 3. A diagonal RNN $f_{\theta}$ consisting of a sequence mixer $\mathbf{\Lambda}_{t}$ and a channel mixer $\mathbf{Q}_{t}$ is iterated until convergence towards the hidden states of an implicitly dense RNN $F_{\theta}$. (b) FP-RNN variants with channel mixer introduced in Sec. 3.3 and 3.4 solve the state-tracking task $A_{5}$ up to various sequence lengths. (c) FP-RNNs adapt their computation time to the difficulty of the task by varying the number fixed-point iterations $\ell^{*}$
3.2 The Fixed-Point Iteration
Solving fixed-point equations such as $h=f_{\theta}(x,h)$, is perhaps one of the most well-studied problems in mathematics (Granas et al., 2003). In the context of deep learning, the literature on Neural ODEs (Chen et al., 2018) and Deep Equilibrium Models (Bai et al., 2019; Ghaoui et al., 2021) investigates fixed-point methods for implicit parametrizations of neural networks. A straightforward, yet effective method computes the forward pass by simply rolling out the fixed-point iteration. In the context of solving $h^{*}=f_{\theta}(x,h^{*})$, this corresponds to introducing an iteration in depth $h^{\ell}=f_{\theta}(x,h^{\ell-1})$. Denoting $\ell$ as the current iteration in depth (i.e., over the layer dimension), and $t$ as the current iteration in time (i.e., over the sequence dimension), the iteration starts at $h_{t}^{0}=0$ and proceeds with
$h_{t}^{\ell}=\mathbf{\Lambda}_{t}h_{t-1}^{{\ell}}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{\ell-1}.$ (6)
Intuitively, this iteration mixes information with interleaved channel mixing (with $\mathbf{Q}_{t}$) and sequence mixing (with $\mathbf{\Lambda}_{t}$) until convergence towards the hidden states of an implicit dense RNN $F_{\theta}$ (cf. 9(a)).
The difficulty with such an iteration in depth and time is that the recurrent dynamics could explode without proper stabilization. While the recurrence in time can be stabilized with RNN techniques  (Zucchet & Orvieto, 2024) such as an input gate $\mathbf{I}-\mathbf{\Lambda}_{t}$, the recurrence in depth, however, could still diverge if $f_{\theta}(x,h)$ does not have an attracting fixed-point (Granas et al., 2003). In order to design a diagonal linear RNN $f_{\theta}(x,h)$ which is guaranteed to have an attracting fixed-point, we make use of Banach (1922)’s theorem. In our context, the theorem states that $f_{\theta}(x,h)$ converges to a fixed-point from any initialization $h^{0}$ if it has a Lipschitz constant $<1$ in $h$. For a fixed-point RNNs with input gate $\mathbf{I}-\mathbf{\Lambda}$, we present the following theorem: